OpenStudy (anonymous):

if y is proportional to 2x-7,and y=-9 when x=2,find x when y=-15.Please show me step by step

OpenStudy (lgbasallote):

proportional which means... y = k(2x-7) -9 = k(2(2) - 7) -9 = k(4 - 7) -9 = k(-3) -9 = -3k divide both sides by -3.... 3 = k when y = -15... -15 = (3)(2x-7) -15 = 6x - 21 21 - 15 = 6x 6 = 6x divide both sides by 6... 1 = x

OpenStudy (lgbasallote):

k is the proportionality constant

OpenStudy (anonymous):

Thank you so much